OpenStudy (anonymous):

Simplify x^2-36/ x^2+3x -18

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x^2-36/ x^2+3x -18\] Multiply by \[(x^2/x^2)\] (which is the same thing as 1) to get: \[(x^4+3x^3-18x^2-36)/(x^2)\]

OpenStudy (anonymous):

Hint: Think of the method for GCD (Greatest Common Denominator) ;-)

OpenStudy (anonymous):

i dont get it :"(

OpenStudy (anonymous):

did you understand mine?

OpenStudy (anonymous):

agentx5 must misunderstood it

OpenStudy (anonymous):

whats the over all answer though

OpenStudy (anonymous):

What two things can be cancel ??

OpenStudy (anonymous):

x+6?

OpenStudy (anonymous):

so its -6/-3?

OpenStudy (anonymous):

is should be (x-6)/(x-3)

OpenStudy (anonymous):

sorry i left the x's out. and THANKS! Deff one of the most helpful people

OpenStudy (anonymous):

did you understand how i factor out the equations from the original problem?

OpenStudy (anonymous):

ehh not really. what hapened with the -18

OpenStudy (anonymous):

Helpful tip: parenthesis are very important! ^_^

OpenStudy (anonymous):

Additional helpful tip: 1. Click the "∑ Equation" button below 2. Type "frac {numerator}{denominator}" (but no quotes, logically, and replace the words with your equation or expression's numerator or denominator) 3. Insert & tada! Looks like this: \[\frac {numerator}{denominator}\]

OpenStudy (anonymous):

\[x^2−36/x^2+3x−18 \neq \frac{x^2−36}{x^2+3x−18}=\frac{(x−6)(x+6)}{(x−3)(x+6)}=\frac{x−6}{x−3}\] What I wrote was correct based on what I saw, @OsmundF \[x^2−36/x^2+3x−18\] \[(\frac{x^2}{1}−\frac{36}{x^2}+\frac{3x}{1}−\frac{18}{1} ) , GCD = x^2\] \[\frac{(x^2*x^2+3 x*x^2-18 *x^2-36*(x^2/x^2))}{x^2}\] \[\frac{(x^4+3 x^3-18 x^2-36)}{x^2}\] ^_^ (and depending on which problem you intended, @Haleybob, now you have answers for each one and processes on how to repeat for similar problems!)

OpenStudy (anonymous):

Oh and to answer your follow question, the -18 = (-3) * (+6). In reverse, take -18 / +6 and see what you get. :-)